Heather is 15 years older than Kim. In 10 years heather will be twice as old as Kim will be. How old is heather now
Answer by stanbon(57307)

Heather is 15 years older than Kim. In 10 years heather will be twice as old as Kim will be. How old is heather now
-----
Equations:
h = k + 15
(h+10) = 2(k+10)
----
Substitute for "h" and solve for "k":
k+15 + 10 = 2k + 20
k + 25 = 2k + 20
k = 5 (Kim's age now)
------
Solve for "h":
h = k + 15
h = 5 + 15 = 20 yrs (Heather's age now)
===================
Cheers,
Stan H.
======================